# Sim2Real in Reconstructive Spectroscopy: Deep Learning with Augmented Device-Informed Data Simulation

### v2

@@ I Introduction
Optical spectroscopy is a versatile technique for various scientific, industrial, and consumer applications. Recently, computational spectroscopy using reconstructive algorithms Chang and Lee (2008); Kurokawa, Choi, and Chang (2011); Bao and Bawendi (2015); Wang and Yu (2014); Wang et al. (2019); Kim et al. (2019); Sarwar et al. (2020); Brown et al. (2021a) has been rapidly developed owing to its potential to enable a miniaturized spectrometer Meng, Cadusch, and Crozier (2021); Zhang et al. (2018a). A spectrometer encodes optical spectral information spatially or temporally and then measures the encoded information using a series of photodetectors. The transformation between the input signal $\bm{x}$ and the photo-detector readout $\bm{y}$ in a spectrometer design can be either linear or nonlinear. However, a linear design is typically preferred. An example of a nonlinear encoding is observed in Fourier-Transform infrared (FTIR) spectrometers, which produce an auto-correlation of the input through a variable time-delay interferometer. For a linear system, $\bm{x}$ and $\bm{y}$ can be represented using column vectors while the mapping is represented by a responsivity matrix $\bm{R}$ such that $\bm{y}=\bm{Rx}$. For example, a monochromator spatially disperses different spectral components with a block-diagonal $\bm{R}$ matrix. In a reconstructive spectrometer, $\bm{R}$ is generally much more complex. Each photo-detector’s readout depends on all instead of a few spectral components. While the complex structure of $\bm{R}$ means a lot of computational resources are generally required to recover the spectrum $\bm{x}$ from the photo-detector readouts $\bm{y}$, it also opens up new opportunities. These include the ability to tailor $\bm{R}$ to encode the spectral information directly to parameters of interest in the application, such as the spectral peak positions and the relative intensities between the peaks; the ability to increase the accuracy of spectral reconstruction by focusing on the most important spectral information even when $\bm{y}$’s dimension is much smaller than $\bm{x}$’sCandes, Romberg, and Tao (2006); and the ability to miniaturize the spectrometer.
@@
A wealth of research has been conducted to increase spectral reconstruction’s efficiency and accuracy. The approaches can generally be categorized as optimization-based or data-driven. The optimization-based approaches formulate the inverse problem into a convex optimization problem in terms of the signal to be reconstructed. To deal with non-uniqueness of the solution, different types of regularizations (e.g., nonnegativity, sparsity, and smoothness) have been considered Rudin, Osher, and Fatemi (1992); Sarwar et al. (2022); Chang and Lee (2008). The effectiveness of these optimization-based techniques often hinges on the precise adjustment of regularization parameters and hyperparameters, a task that can be challenging Li et al. (2023). Moreover, optimization-based approaches based on iterative procedures can be energy and computationally expensive for critical applications such as Internet-of-Things or time-sensitive tasks. Despite the challenges, miniaturized spectrometers based on these spectral-reconstruction techniques have reported good performance. For example, we recently developed a chip-scale spectrometer concept by integrating 16 semiconductor photo-detectors, each with a different spectral response Sarwar et al. (2022). We designed the photo-detectors using nanostructured materials to minimize the dependence on the incident angle of light and thus eliminated the need for any external optical components. The entire spectrometer, sans the electronic readout circuitry, is only a few microns thick. As the photodetector’s response was broadband, the device was able to recover basic spectral information, including the positions and relative intensities of peaks across the entire visible spectrum. The accuracy of reconstructing a multi-peak spectrum were 0.97% (RMS error) for locating the peak locations, comparable to a monochromator with twice the number of detectors. Enabled by the rich interaction between different spectral components in each photo-detector, we also showed that only seven detectors were sufficient for input signals with up to three peaks and had relatively smooth spectra.
@@
Figure 1: The diagram of our proof-of-concept Sim2Real framework in the reconstructive spectroscopy. The orange arrows denote existing methods, which require collecting and training on real-world data. The blue box and arrows denote our proposed Sim2Real framework that effectively address the domain shift between simulated and real-world data. The domain shift is visualized through PCA in Figure 2 See Section II for details. The Response Matrix plot is reprinted with permission from [13]. Copyright 2022 American Chemical Society.
@@
In comparison, deep learning-based data-driven approaches Brown et al. (2021a); Kim, Park, and Lee (2020); Zhang, Zhu, and Bao (2020a) showed the potential to address the challenges of an optimization-based approach. First, deep overparametrized networks trained by gradient descent methods tend to enforce the regularization implicitlyArora et al. (2019), avoiding the need to adjust regularization parameters. Second, while training neural networks can be computationally expensive, reconstructing a new measurement during inference requires only a single forward pass through the network. In contrast, optimization-based approaches can be costly in terms of energy and computing resources during inference. Various data-driven methods of reconstructive spectroscopy have been developed, including utilizing a simple, fully connected network on a plasmonic spectral encoder on the deviceBrown et al. (2021a), minimizing potential noise before the reconstruction process on a colloidal quantum dot spectrometerZhang, Zhu, and Bao (2020a), and a specifically tailored residual convolutional neural network aiming to improve the reconstruction performanceKim, Park, and Lee (2020).
@@
A significant challenge with the current deep learning methodologies is their dependence on extensive volumes of real, precisely annotated training data to achieve optimal performanceBrown et al. (2021a); Li et al. (2023). In a practical setting, gathering such real, labeled data for spectrometer reconstruction is both costly and time-intensive. Additionally, the actual labeled data pairs collected often contain considerable noise, which may result in poor performance during testing when used to train deep learning models.
@@
The rest of the paper is organized as follows. In Section II, we introduce the mathematical formulation of the spectral reconstruction problem while identifying the limitations of an existing method. In Section III, we present our main contributions and discuss how they address the limitations of the existing method. In Section IV, we validate the performance of our proposed method on real-world data by comparing it with the state-of-the-art method and discussing the implications.
II Problem formulation and existing Methods
@@
II.1 Spectral Reconstruction Problem
@@
II.2 Optimization-Based Approaches and Limitations
Note that given the response matrix $\bm{R}$, the reconstruction problem in (3) can be approached using (non-negative) least squares, as discussed in the introduction. Due to the ill-posedness of the problem, explicit regularization techniques are used to ensure unique solutions with desirable additional structures. Within optimization-based methods, the least-squares (LS) estimate $\hat{\bm{x}}\in\mathrm{argmin}_{\bm{x}}\|\bm{y}-\bm{Rx}\|_{2}^{2}$ fails to consider the signal’s nonnegativity, often leading to suboptimal solutions Kurokawa, Choi, and Chang (2010). This issue can be addressed by employing nonnegative least squares (NNLS) Chang and Lee (2008). To deal with an underdetermined system, regularization strategies like $\ell_{1}$-norm, $\ell_{2}$-norm, or total-variation (TV) norm regularization have also been proven effectiveRudin, Osher, and Fatemi (1992); Sarwar et al. (2022).
@@ II.2 Optimization-Based Approaches and Limitations
II.3 Deep Learning Methods and Challenges
@@
Therefore, based on the problem formulation in Equation 3, previous approaches have focused on using the response matrix $\bm{R}$ to produce simulated spectral data pairs $(\bm{y}_{\textsf{sim}},\bm{x}_{\textsf{sim}})$, aiming to mimic the distribution of real data obtained in laboratory settings. However, we observe that the simulated data and the real-measured data exhibit an undesirable phenomenon known as “domain shift”, leading to the poor performance of model trained solely on simulated data when applied to real ones. To provide a straightforward quantitative assessment of the domain shift in our problem, here we ran PCA on both simulated data and real data (preprocessed by log-min-max as mentioned in Section IV) to visualize the discrepancy via their two principal components. As depicted in Figure 2, the real data presents a much larger spread and variation along both axes compared to the simulated data, which demonstrated the existence of the domain shift.
@@
Later in Section III, we delve into this issue and detail our approach to bridging the domain shift. For reader’s convenience, we first review previous approaches in detail.
@@
In practice, the simulated spectral signals $\bm{x}_{\textsf{sim}}$ can be viewed as the sum of Lorentzian distribution functions Kim, Park, and Lee (2020); Sarwar et al. (2022). Given the response matrix $\bm{R}$, a simulated spectral signal $\bm{x}_{\textsf{sim}}$ and its corresponding encoded signal $\bm{y}_{\textsf{sim}}$ can be generated as follows:
@@
We illustrate this in Figure 4, where we train a model exclusively on simulated data that can successfully reconstruct the spectrum given simulated data as input. However, the model’s performance degrades significantly on real-world data, even when the real-world and simulated data appear visually similar. Detailed empirical evidences are discussed in Section IV.4.
@@
Figure 4: The testing RMSE on the simulated and real dataSarwar et al. (2022) for ResCNN Kim, Park, and Lee (2020) and our proposed model ReSpecNN. Both models followed the Sim2Real training setting, that is, trained solely on simulated data. Our model further incorporated the hierarchical data augmentation during training, while ResCNN does not.
II.4 Domain Shift between “Sim” and “‘Real”
@@
The challenges of “domain shift” are common for deep learning-based approaches. The “domain shift” or “domain gap” issue was first observed in transfer learningPan and Yang (2009), where a model trained on a large source dataset is fine-tuned on a smaller, more specific dataset. To mitigate the performance degradation due to distributional discrepancy between source and target datasets, many techniques have been proposed, including domain adaptationTzeng et al. (2017); Ganin and Lempitsky (2015); Zhang et al. (2018b), meta-learningFinn, Abbeel, and Levine (2017), and few-shot learningVinyals et al. (2016); Snell, Swersky, and Zemel (2017). These techniques have been shown to be effective in reducing the domain gap in various fields including computer visionGirshick et al. (2014); Zoph et al. (2018) and natural language processingDai and Le (2015); Howard and Ruder (2018); Devlin et al. (2019).
@@
However, when applied to the spectroscopy reconstruction, those proposed methods fall short of solving the domain gap for the following reasons. First, we do not assume access to any data from the target distribution. Existing methods still require a certain amount of data from the targeted domain for fine-tuning, which is not applicable under our assumption. Second, most domain adaptation methods have been designed for classification-based tasks, making them less directly applicable to our spectral reconstruction problem in spectroscopy, which is fundamentally a (non-negative) regression problem. While a few existing results focus on practical regression problems Chen et al. (2021); Nejjar, Wang, and Fink (2023), they often deal with simple regression problems rather than tackling the more challenging inverse problem setting which we consider. Third, regarding the topic of deep learning for inverse problems, for tasks like image reconstruction, existing research primarily leverages the inductive bias of neural network architectures suitable for image signals Dong et al. (2015); Kulkarni et al. (2016); Ongie et al. (2020). In this work on spectral reconstruction, we deal with one-dimensional spectral signals which requires a different model design.
@@
While deep learning approaches have become commonly used for spectral reconstruction problems, few studies tackle such domain shift issues directly. A recent workBrown et al. (2021a) successfully reconstructed spectra with up to 14 peaks using a model trained on spectra with up to 8 peaks under the blind testing conditions, the training and test data are still drawn from similar conditions and distributions. Therefore, a specialized method is required to tackle the domain shift issue in solving the spectral reconstruction problem. We now introduce the proposed Sim2Real framework in the following.
III Our Sim2Real Framework
@@
III.1 Hierarchical Data Augmentation
Although we measure the response matrix $\bm{R}$ in advance and consider it to be fixed and known, our hierarchical data augmentation strategy acknowledges the potential uncertainty in our measurement of $\bm{R}$ and the encoded signal vector $\bm{y}$ to improve model robustness and minimize the domain gap. For every pair of simulated training data $(\bm{y}_{\textsf{sim}},\bm{x}_{\textsf{sim}})$ introduced in Section II.3, we systematically introduce noise as outlined in Algorithm 1.
@@ III.1 Hierarchical Data Augmentation
   $\bm{x}_{\textsf{sim}}$: a simulated spectral signal from Section II.3
@@
To perturb the response matrix $\bm{R}$, we simply add Gaussian noise with zero mean and variance related to entries of $\bm{R}$. That is, for each noisy perturbation matrix $\bm{\Delta}_{s}$ ($s=1,\cdots,S$), its $(i,j)^{th}$ entry is sampled i.i.d from the distribution $\mathcal{N}\left(0,\ \sigma^{2}_{ij}\right)$ with $\sigma_{ij}=\alpha\cdot R_{ij}$, where $\alpha$ is a hyperparameter controlling the intensity of perturbation on $\bm{R}$. For perturbing the encoded signal $\bm{y}$, we inject nonnegative noise $\bm{\epsilon}_{t}$ ($t=1,\cdots,T$). Here, each entry of $\bm{\epsilon}_{t}$ is sampled i.i.d from the Gaussian distribution and then passed through the $\textsf{ReLU}(z)=\max\{0,z\}$ operator to enforce non-negativity. In practice, we determine those hyper-parameters through the empirical experiments. Details could be found in the later section IV.1.
@@
III.2 Network Architecture: ReSpecNN

### v3

@@ 1 Introduction
Optical spectroscopy is a versatile technique for various scientific, industrial, and consumer applications. Recently, computational spectroscopy using reconstructive algorithms [1, 2, 3, 4, 5, 6, 7, 8] has been rapidly developed owing to its potential to enable a miniaturized spectrometer [9, 10]. A spectrometer encodes optical spectral information spatially or temporally and then measures the encoded information using a series of photodetectors. The transformation between the input signal $\bm{x}$ and the photo-detector readout $\bm{y}$ in a spectrometer design can be either linear or nonlinear. However, a linear design is typically preferred. An example of a nonlinear encoding is observed in Fourier-Transform infrared (FTIR) spectrometers, which produce an auto-correlation of the input through a variable time-delay interferometer. For a linear system, $\bm{x}$ and $\bm{y}$ can be represented using column vectors while the mapping is represented by a responsivity matrix $\bm{R}$ such that $\bm{y}=\bm{Rx}$. For example, a monochromator spatially disperses different spectral components with a block-diagonal $\bm{R}$ matrix. In a reconstructive spectrometer, $\bm{R}$ is generally much more complex. Each photo-detector’s readout depends on all instead of a few spectral components. While the complex structure of $\bm{R}$ means a lot of computational resources are generally required to recover the spectrum $\bm{x}$ from the photo-detector readouts $\bm{y}$, it also opens up new opportunities. These include the ability to tailor $\bm{R}$ to encode the spectral information directly to parameters of interest in the application, such as the spectral peak positions and the relative intensities between the peaks; the ability to increase the accuracy of spectral reconstruction by focusing on the most important spectral information even when $\bm{y}$’s dimension is much smaller than $\bm{x}$’s[11]; and the ability to miniaturize the spectrometer.
@@
A wealth of research has been conducted to increase spectral reconstruction’s efficiency and accuracy. The approaches can generally be categorized as optimization-based or data-driven. The optimization-based approaches formulate the inverse problem into a convex optimization problem in terms of the signal to be reconstructed. To deal with non-uniqueness of the solution, different types of regularizations (e.g., nonnegativity, sparsity, and smoothness) have been considered [12, 13, 1]. The effectiveness of these optimization-based techniques often hinges on the precise adjustment of regularization parameters and hyperparameters, a task that can be challenging [14]. Moreover, optimization-based approaches based on iterative procedures can be energy and computationally expensive for critical applications such as Internet-of-Things or time-sensitive tasks. Despite the challenges, miniaturized spectrometers based on these spectral-reconstruction techniques have reported good performance. For example, we recently developed a chip-scale spectrometer concept by integrating 16 semiconductor photo-detectors, each with a different spectral response [13]. We designed the photo-detectors using nanostructured materials to minimize the dependence on the incident angle of light and thus eliminated the need for any external optical components. The entire spectrometer, sans the electronic readout circuitry, is only a few microns thick. As the photodetector’s response was broadband, the device was able to recover basic spectral information, including the positions and relative intensities of peaks across the entire visible spectrum. The accuracy of reconstructing a multi-peak spectrum were 0.97% (RMS error) for locating the peak locations, comparable to a monochromator with twice the number of detectors. Enabled by the rich interaction between different spectral components in each photo-detector, we also showed that only seven detectors were sufficient for input signals with up to three peaks and had relatively smooth spectra.
@@
Figure 1: The diagram of our proof-of-concept Sim2Real framework in the reconstructive spectroscopy. The orange arrows denote existing methods, which require collecting and training on real-world data. The blue box and arrows denote our proposed Sim2Real framework that effectively address the domain shift between simulated and real-world data. The domain shift is visualized through PCA in Figure 2 See Section 2 for details. The Response Matrix plot is reprinted with permission from [13]. Copyright 2022 American Chemical Society.
@@
In comparison, deep learning-based data-driven approaches [8, 15, 16] showed the potential to address the challenges of an optimization-based approach. First, deep overparametrized networks trained by gradient descent methods tend to enforce the regularization implicitly[17], avoiding the need to adjust regularization parameters. Second, while training neural networks can be computationally expensive, reconstructing a new measurement during inference requires only a single forward pass through the network. In contrast, optimization-based approaches can be costly in terms of energy and computing resources during inference. Various data-driven methods of reconstructive spectroscopy have been developed, including utilizing a simple, fully connected network on a plasmonic spectral encoder on the device[8], minimizing potential noise before the reconstruction process on a colloidal quantum dot spectrometer[16], and a specifically tailored residual convolutional neural network aiming to improve the reconstruction performance[15].
@@
A significant challenge with the current deep learning methodologies is their dependence on extensive volumes of real, precisely annotated training data to achieve optimal performance[8, 14]. In a practical setting, gathering such real, labeled data for spectrometer reconstruction is both costly and time-intensive. Additionally, the actual labeled data pairs collected often contain considerable noise, which may result in poor performance during testing when used to train deep learning models.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the mathematical formulation of the spectral reconstruction problem while identifying the limitations of an existing method. In Section 3, we present our main contributions and discuss how they address the limitations of the existing method. In Section 4, we validate the performance of our proposed method on real-world data by comparing it with the state-of-the-art method and discussing the implications.
2 Problem formulation and existing Methods
@@
2.1 Spectral Reconstruction Problem
@@
2.2 Optimization-Based Approaches and Limitations
Note that given the response matrix $\bm{R}$, the reconstruction problem in (3) can be approached using (non-negative) least squares, as discussed in the introduction. Due to the ill-posedness of the problem, explicit regularization techniques are used to ensure unique solutions with desirable additional structures. Within optimization-based methods, the least-squares (LS) estimate $\hat{\bm{x}}\in\mathrm{argmin}_{\bm{x}}\|\bm{y}-\bm{Rx}\|_{2}^{2}$ fails to consider the signal’s nonnegativity, often leading to suboptimal solutions. This issue can be addressed by employing nonnegative least squares (NNLS) [1]. To deal with an underdetermined system, regularization strategies like $\ell_{1}$-norm, $\ell_{2}$-norm, or total-variation (TV) norm regularization have also been proven effective[12, 13].
@@ 2.2 Optimization-Based Approaches and Limitations
2.3 Deep Learning Methods and Challenges
@@
Therefore, based on the problem formulation in Equation 3, previous approaches have focused on using the response matrix $\bm{R}$ to produce simulated spectral data pairs $(\bm{y}_{\textsf{sim}},\bm{x}_{\textsf{sim}})$, aiming to mimic the distribution of real data obtained in laboratory settings. However, we observe that the simulated data and the real-measured data exhibit an undesirable phenomenon known as “domain shift”, leading to the poor performance of model trained solely on simulated data when applied to real ones. To provide a straightforward quantitative assessment of the domain shift in our problem, here we ran PCA on both simulated data and real data (preprocessed by log-min-max as mentioned in Section 4) to visualize the discrepancy via their two principal components. As depicted in Figure 2, the real data presents a much larger spread and variation along both axes compared to the simulated data, which demonstrated the existence of the domain shift.
@@
Later in Section 3, we delve into this issue and detail our approach to bridging the domain shift. For reader’s convenience, we first review previous approaches in detail.
@@
In practice, the simulated spectral signals $\bm{x}_{\textsf{sim}}$ can be viewed as the sum of Lorentzian distribution functions [15, 13]. Given the response matrix $\bm{R}$, a simulated spectral signal $\bm{x}_{\textsf{sim}}$ and its corresponding encoded signal $\bm{y}_{\textsf{sim}}$ can be generated as follows:
@@
We illustrate this in Figure 3, where we train a model exclusively on simulated data that can successfully reconstruct the spectrum given simulated data as input. However, the model’s performance degrades significantly on real-world data, even when the real-world and simulated data appear visually similar. Detailed empirical evidences are discussed in Section 4.4.
@@
Figure 3: The testing RMSE on the simulated and real data[13] for ResCNN [15] and our proposed model ReSpecNN. Both models followed the Sim2Real training setting, that is, trained solely on simulated data. Our model further incorporated the hierarchical data augmentation during training, while ResCNN does not.
2.4 Domain Shift between “Sim” and “‘Real”
@@
The challenges of “domain shift” are common for deep learning-based approaches. The “domain shift” or “domain gap” issue was first observed in transfer learning[18], where a model trained on a large source dataset is fine-tuned on a smaller, more specific dataset. To mitigate the performance degradation due to distributional discrepancy between source and target datasets, many techniques have been proposed, including domain adaptation[19, 20, 21], meta-learning[22], and few-shot learning[23, 24]. These techniques have been shown to be effective in reducing the domain gap in various fields including computer vision[25, 26] and natural language processing[27, 28, 29].
@@
However, when applied to the spectroscopy reconstruction, those proposed methods fall short of solving the domain gap for the following reasons. First, we do not assume access to any data from the target distribution. Existing methods still require a certain amount of data from the targeted domain for fine-tuning, which is not applicable under our assumption. Second, most domain adaptation methods have been designed for classification-based tasks, making them less directly applicable to our spectral reconstruction problem in spectroscopy, which is fundamentally a (non-negative) regression problem. While a few existing results focus on practical regression problems [30, 31], they often deal with simple regression problems rather than tackling the more challenging inverse problem setting which we consider. Third, regarding the topic of deep learning for inverse problems, for tasks like image reconstruction, existing research primarily leverages the inductive bias of neural network architectures suitable for image signals [32, 33, 34]. In this work on spectral reconstruction, we deal with one-dimensional spectral signals which requires a different model design.
@@
While deep learning approaches have become commonly used for spectral reconstruction problems, few studies tackle such domain shift issues directly. A recent work[8] successfully reconstructed spectra with up to 14 peaks using a model trained on spectra with up to 8 peaks under the blind testing conditions, the training and test data are still drawn from similar conditions and distributions. Therefore, a specialized method is required to tackle the domain shift issue in solving the spectral reconstruction problem. We now introduce the proposed Sim2Real framework in the following.
3 Our Sim2Real Framework
@@
3.1 Hierarchical Data Augmentation
Although we measure the response matrix $\bm{R}$ in advance and consider it to be fixed and known, our hierarchical data augmentation strategy acknowledges the potential uncertainty in our measurement of $\bm{R}$ and the encoded signal vector $\bm{y}$ to improve model robustness and minimize the domain gap. For every pair of simulated training data $(\bm{y}_{\textsf{sim}},\bm{x}_{\textsf{sim}})$ introduced in Section 2.3, we systematically introduce noise as outlined in Algorithm 1.
@@ 3.1 Hierarchical Data Augmentation
    $\bm{x}_{\textsf{sim}}$: a simulated spectral signal from Section 2.3
@@
Algorithm 1 Hierarchical Data Augmentation (HDA)
To perturb the response matrix $\bm{R}$, we simply add Gaussian noise with zero mean and variance related to entries of $\bm{R}$. That is, for each noisy perturbation matrix $\bm{\Delta}_{s}$ ($s=1,\cdots,S$), its $(i,j)^{th}$ entry is sampled i.i.d from the distribution $\mathcal{N}\left(0,\ \sigma^{2}_{ij}\right)$ with $\sigma_{ij}=\alpha\cdot R_{ij}$, where $\alpha$ is a hyperparameter controlling the intensity of perturbation on $\bm{R}$. For perturbing the encoded signal $\bm{y}$, we inject nonnegative noise $\bm{\epsilon}_{t}$ ($t=1,\cdots,T$). Here, each entry of $\bm{\epsilon}_{t}$ is sampled i.i.d from the Gaussian distribution and then passed through the $\textsf{ReLU}(z)=\max\{0,z\}$ operator to enforce non-negativity. In practice, we determine those hyper-parameters through the empirical experiments. Details could be found in the later section 4.1.
@@
3.2 Network Architecture: ReSpecNN